(1). Bertha is three times as old as Martha but four times as old as Gifty. Fine their respective ages if the sum of their ages is 84 years.
Answer by stanbon(75887)

Bertha is three times as old as Martha but four times as old as Gifty. Find their respective ages if the sum of their ages is 84 years.
-----
Equations:
B = 3*M
B = 4*G
B + M + G = 84
-------------------------
B + (1/3)B + (1/4)B = 84
------------
12B + 4B + 3B = 12*84
--------------------------------
19B = 12*84
B = 53.05 years
----
M = (1/3)53.05 = 17.68 years
B = (1/4)53.05 = 13.26 years
------------
Cheers,
Stan H.
------------